Suppose that a firms demand curve is given by P 14 0.5 Q What is the profit-maximizing price if total cost is TC 3.Q?
Answer #1

Answer

the firm produces at MR=MC

MC=dTC/dQ=3

MR=14-Q ........ an MR curve is double sloped than an inverse linear demand curve

equating both

14-Q=3

Q=11 units

the firm produces 11 units

P=14-0.5*11

=8.5

The price is $8.5